### v3

@@ 1 Introduction
Figure 1: MemOS achieves state-of-the-art performance across all reasoning tasks. This figure summarizes LLM-Judge scores on the LOCOMO benchmark, covering four task categories (Single-hop, Multi-hop, Open-domain, Temporal Reasoning) and the overall average. MemOS (MemOS-0630) consistently ranks first in all categories, outperforming strong baselines such as mem0, LangMem, Zep, and OpenAI-Memory, with especially large margins in challenging settings like multi-hop and temporal reasoning. Error bars indicate standard deviation. Full metric breakdown is provided in Table 3.
@@ 6 Evaluation
To systematically evaluate the capabilities of MemOS, we conduct both holistic and component-level experiments. We begin by benchmarking the full system on the LOCOMO benchmark suite to assess its performance in memory-intensive reasoning tasks, comparing against several state-of-the-art baselines. In addition, we present targeted evaluations of key architectural subsystems, including multi-perspective memory organization, hybrid semantic retrieval, task-aligned scheduling, and KV-based activation memory injection. These experiments assess the individual effectiveness of each component and its contribution to overall system performance.
6.1 End-to-End Evaluation on LOCOMO
Table 3: Comparison of LLM Judge Scores across five major tasks in the LOCOMO benchmark. Each bar shows the mean evaluation score judged by LLMs for a given method-task pair, with standard deviation as error bars. MemOS-0630 consistently outperforms baseline methods (LangMem, Zep, OpenAI, Mem0) across all task types, especially in multi-hop and temporal reasoning scenarios.
Category Method
Chunk /
Mem Tok
Top-K LLMJudge Score F1 RL B1 B2 METEOR BERT-F1 Sim
@@ 6.1 End-to-End Evaluation on LOCOMO
zep 2320 20 50.42±0.29 32.49 35.07 27.38 18.81 28.49 35.26 66.95
openai 4141 - 61.83±0.10 36.96 41.45 30.72 22.02 35.39 40.56 69.84
@@
[…]
zep 2351 20 42.20±0.77 23.14 24.63 14.96 8.49 17.38 25.15 64.26
[…]
zep 2276 20 38.19±0.49 19.76 20.62 13.17 6.18 14.20 21.07 58.59
openai 4071 - 32.99±1.30 17.19 15.88 11.04 5.23 12.05 19.37 57.53
[…]
zep 2295 20 19.11±0.29 17.59 19.03 14.57 8.11 13.81 17.59 59.38
openai 4048 - 28.25±0.59 23.90 24.47 18.25 11.87 19.35 23.11 59.53
[…]
temporal reasoning memos-0630 1655 20 73.21±0.25 53.67 53.69 46.37 29.69 43.45 48.48 76.97
langmem 165 - 55.76±0.16 39.18 41.01 32.59 21.27 32.28 42.03 74.76
zep 2318 20 41.62±0.21 26.88 28.91 21.55 13.90 22.51 28.84 64.36
openai 4077 - 52.75±0.08 32.30 35.20 25.64 17.64 29.10 34.10 66.74
[…]
overall memos-0630 1593 20 73.31±0.05 44.42 47.65 36.88 25.43 40.20 44.15 73.51
We evaluate MemOS against a diverse set of strong baselines, each representing a distinct memory system design paradigm. Specifically, LangMem applies hierarchical semantic retrieval over flat textual history; Zep integrates time-aware knowledge graphs with structured query resolution; OpenAI-Memory represents a commercial, closed-source memory module with opaque internal logic; and Mem0 implements slot-based long-term memory with top-k semantic search. To ensure architectural parity, all methods are implemented over the same LLM backbone (GPT-4o-mini).
All experiments are conducted on an 80GB H800 GPU under identical hardware and software configurations. For memory-augmented systems, we vary the number of retrieved items (Top-K) and chunk granularity (Chunk / Mem Tok), which controls the length of each retrieved memory segment. The configuration for each method is selected based on its best validation performance, ensuring a fair and optimized comparison across all metrics.
We report LLM-judge scores as the primary evaluation metric (Table 3), supported by standard generation quality indicators including F1, ROUGE-L (RL), BLEU-1/2 (B1/B2), METEOR, and BERTScore-F1 (BERT-F1), as well as cosine similarity (Sim) computed over semantic embeddings.
Overall, MemOS achieves the best average performance across all task categories, consistently outperforming strong baselines such as mem0, openai-memory, and zep. Across all sub-tasks in the LOCOMO benchmark, MemOS ranks among the top performers, maintaining first or second place in nearly every category. It demonstrates clear advantages in multi-hop and temporal reasoning, where long-range memory and contextual integration are especially critical. Beyond LLM-judge scores, MemOS also delivers strong generation quality across F1, ROUGE-L, and BLEU, particularly in long-form completeness and stylistic consistency. At the representation level, it maintains tight semantic alignment with reference answers, as indicated by consistently high cosine similarity in semantic embeddings across tasks.
@@
Table 4: Latency and LLM evaluation scores across various methods on the LOCOMO benchmark. RAG is evaluated under different chunk sizes and retrieval depths (Top-K = 1 or 2), while other baselines include standard retrieval systems and memory-augmented models. Metrics include LLMJudge scores (evaluating answer quality), search latency (P50/P95), and total end-to-end latency (P50/P95). MemOS-0630 achieves the highest LLM score with competitive latency performance.
Method
Chunk /
Mem Tok
Top-K LLMJudge Scores search duration (ms) total duration (ms)
P50 P95 P50 P95
1 44.61±0.05 516 800 1306 1963
128 2 55.71±0.05 523 850 1325 2040
1 45.13±0.19 553 1288 1438 2606
256 2 56.54±0.25 575 1371 1496 2843
1 41.36±0.24 481 1979 1331 4129
512 2 54.29±0.19 482 2070 1351 4252
1 36.04±0.09 1008 2436 2061 4443
1024 2 46.97±0.03 468 808 1466 2193
1 33.70±0.05 460 986 1387 2311
2048 2 44.81±0.05 456 903 1476 2479
1 33.9±0.05 449 715 1432 2324
4096 2 48.53±0.13 459 1055 1606 3324
1 41.45±0.17 692 1733 2016 5037
RAG 8192 2 58.20±0.08 688 1773 2335 6008
Full-Context 22636 - 71.58±0.08 - - 2339 7016
langmem 165 - 55.76±0.16 17226 29344 18025 30139
zep 2318 20 41.62±0.21 1364 1901 9777 20197
openai 4077 - 52.75±0.08 - - 1184 2240
mem0 1171 20 64.57±0.06 1297 1416 4906 5962
memos-0630 1593 20 73.31±0.05 1758 1969 4942 7937
6.2 Evaluation of Memory Retrieval
We conduct a focused evaluation to analyze the efficiency and effectiveness of memory retrieval across representative system designs. As shown in Table 4, we compare latency and generation quality under different memory configurations, including standard RAG pipelines, memory-augmented models, and our proposed MemOS.
To test RAG-style retrieval systems, we systematically vary chunk sizes (from 128 to 8192 tokens) and Top-K values (1 or 2) to observe the trade-offs between context size, search latency, and LLM output quality. Larger chunk sizes reduce retrieval depth but increase encoding and integration cost, while smaller chunks allow finer granularity at the expense of retrieval breadth.
In addition to standard retrieval, we include full-context and commercial memory systems to establish upper and lower bounds. Notably, the full-context baseline—where the entire dialogue history is loaded into the model—achieves strong LLMJudge scores but suffers from prohibitively high latency due to extreme context length. LangMem and Zep incur substantial retrieval delays from graph traversal or multi-level indexing. OpenAI-Memory offers low latency but only moderate output quality, likely limited by opaque memory heuristics.
Remarkably, MemOS not only matches but surpasses the full-context baseline in LLMJudge scores—while operating at significantly lower latency. Despite managing over 1500 memory tokens, its retrieval time remains close to smaller baselines such as mem0. This demonstrates that MemOS’s hybrid semantic organization and activation-based memory loading can achieve superior performance without the cost of full-context inference.

### v4

@@ 1 Introduction
Figure 1: MemOS achieves state-of-the-art performance across all benchmarks. This figure provides a comprehensive summary of our evaluation results, illustrating the personalized response rate on PreFEval(0 turns and 10 turns), the precision score on PersonaMem, the overall mean score on LongMemEval, and the overall mean LLM judge score on the LoCoMo benchmark. The MemOS (MemOS-1031) consistently ranks first in all categories, significantly outperforming strong baselines such as MIRIX, Mem0, Zep, Memobase, MemU, and Supermemory. Full detailed metric breakdowns are provided in Table 3, Table 5, Table 4 and Table 6.
@@ 6 Evaluation
To systematically evaluate the capabilities of MemOS, we conduct both holistic and component-level experiments. We begin by benchmarking the full system on the LoCoMo [100], LongMemEval [101], PreFEval [102] and PersonaMem [103] benchmark suite to assess its performance in memory-intensive reasoning and personalization tasks, comparing against several state-of-the-art baselines. In addition, we present targeted evaluations of key architectural subsystems, including multi-perspective memory organization, hybrid semantic retrieval, task-aligned scheduling, and KV-based activation memory injection. These experiments assess the individual effectiveness of each component and its contribution to overall system performance.
6.1 End-to-End Evaluation on Long Context Memory
To evaluate long-context and multi-session capabilities, we test MemOS on the LoCoMo [100] and LongMemEval [101] benchmarks against a diverse set of strong baselines, each representing a distinct memory design paradigm.
Specifically, MIRIX [104] manages memory via six specialized components (Core, Episodic, Semantic, Procedural, Resource, and Knowledge Vault); Mem0 [105] implements slot-based long-term memory with top-k semantic search; Zep [106] integrates time-aware knowledge graphs with structured query resolution; Memobase prioritizes the balance between performance, cost, and latency in long-term user memory; Supermemory employs a Dynamic Knowledge Graph to map relationships between memories, enabling semantic understanding as information evolves; and MemU processes multi-modal inputs by extracting and summarizing them into structured memory files. To ensure architectural parity, all methods are implemented over the same LLM backbone (GPT-4o-mini [17]).
All experiments are conducted on an 80GB H800 GPU under identical hardware and software configurations. To ensure a fair and optimized comparison, the configuration for each method is selected based on its best validation performance. Note that for MIRIX, token consumption data is omitted from the results due to system limitations.
Table 3: We evaluated our MemOS-1031 model against various baselines on the LoCoMo benchmark, all using the same GPT-4o-mini foundation LLM. We report the performance based on LLM Judge Scores across five major tasks (single-hop, multi-hop, temporal reasoning, open-domain, and overall), alongside the overall F1 score and the context tokens used by each method. The results demonstrate that MemOS-1031 consistently outperformed baseline methods in almost all scenarios, all while maintaining acceptable context length and a reasonable F1 score.
Method Tokens Single-hop $\uparrow$ Multi-hop $\uparrow$ Temporal Reasoning $\uparrow$ Open-domain $\uparrow$ Overall $\uparrow$ Overall F1 $\uparrow$
MIRIX - 68.22 54.26 68.54 46.88 64.33 28.10
Mem0 1172 73.33 58.75 52.34 45.83 64.57 43.46
Zep 2701 66.23 52.12 54.82 33.33 59.22 41.23
Memobase 2102 73.12 64.65 81.20 53.12 72.01 50.18
MemU 617 66.34 63.12 27.10 50.01 56.55 35.15
Supermemory 500 67.30 51.12 31.77 42.67 55.34 34.87
MemOS-1031 1589 81.09 67.49 75.18 55.90 75.80 45.27
As shown in Table 3, MemOS achieves the best average performance across nearly all task categories on the LoCoMo benchmark. Across all sub-tasks in LoCoMo, MemOS consistently ranks among the top performers, maintaining first or second place in every category. It demonstrates clear advantages in single-hop and multi-hop tasks, where long-range memory and contextual integration are especially critical. Beyond LLM-judge scores, MemOS also delivers strong generation quality in terms of F1 score while maintaining reasonable context length control, implying that its high LLM-judge scores are not a result of retrieval token overflow.
Table 4: We evaluated our MemOS-1031 model against various baselines on the LongMemEval benchmark, all using the same GPT-4o-mini foundation LLM. We report the performance percentage across six key scenarios (single-session preference, single-session assistant, temporal reasoning, multi-session, knowledge update, and single-session user), along with the context tokens used by each method. The results demonstrate that MemOS-1031 achieved the best precision in almost all scenarios while maintaining acceptable control over context length.
@@ 6.1 End-to-End Evaluation on Long Context Memory
MIRIX - 53.3 63.6 25.6 30.1 52.6 72.9 43.49
Zep 1.6k 53.3 75.0 54.1 47.4 74.4 92.9 63.8
Mem0 1.1k 90.0 26.8 72.2 63.2 66.7 82.9 66.4
Memobase 1.5k 80.1 23.2 75.9 66.9 89.7 92.9 72.4
Supermemory 0.4k 89.9 58.9 44.4 52.6 55.1 85.7 58.4
MemU 0.5k 76.7 19.6 17.3 42.1 41.0 67.1 38.4
MemOS-1031 1.4k 96.7 67.9 77.4 70.7 74.3 95.7 77.8
Similarly, for LongMemEval (Table 4), MemOS achieves the best average performance across almost all task categories. Across all sub-tasks in the benchmark, MemOS ranks among the top performers, placing first or second in every category except knowledge updates, and securing the best overall performance on average.
6.2 End-to-End Evaluation on Personalization and Preference Understanding
To evaluate personalization and preference understanding, we test MemOS on the PreFEval [102] and PersonaMem [103] benchmarks against the same baselines detailed in Section 6.1. We use the identical hardware and software configurations as described in Section 6.1. For a fair and optimized comparison, the configuration for each method is selected based on its best validation performance.
Table 5: We evaluated our MemOS-1031 model against various baselines on the PreFEval benchmark [102], all built upon the same GPT-4o-mini[17] foundation LLM. Our comparative analysis involved two scenarios: one without and one with the injection of 10 irrelevant conversation turns. We report the Personalized Response percentage (higher is better) and compare it against the percentages of various error types including Preference Unaware, Preference Hallucination, Inconsistency, and Unhelpful Response (lower is better)—along with the context tokens used by each method. As the results show, MemOS-1031 not only achieved the best Personalized Response performance in both scenarios but also recorded the lowest Preference Unaware error, all while maintaining acceptable error and context length control.
@@ 6.2 End-to-End Evaluation on Personalization and Preference Understanding
with 0 turns Bare LLM 2027.5 80.9 9.1 0.4 0.0 9.6
Bare LLM(+rag-5) 393.7 19.6 25.9 3.3 0.0 51.2
MIRIX - 49.2 9.5 0.0 3.6 37.7
Mem0 83.0 14.0 18.4 1.4 0.3 65.9
Zep 581.3 38.1 18.1 1.7 1.4 40.7
Memobase 430.0 36.0 25.8 1.9 0.3 36.0
Supermemory 117.0 18.5 19.9 2.6 0.6 58.4
MemU 108.1 22.1 20.4 2.2 1.1 54.2
MemOS-1031 557.0 4.6 14.5 1.6 2.1 77.2
with 10 turns Bare LLM 11007.5 93.2 3.9 0.1 0.0 2.8
Bare LLM(+rag-5) 392.5 26.6 27.1 3.9 0.0 43.2
MIRIX - 77.9 72.0 0.0 7.0 7.9
Mem0 90.0 14.8 18.4 3.1 0.0 63.7
Zep 901.2 41.0 15.7 2.1 1.3 39.9
Memobase 563.0 37.0 25.8 2.0 0.1 34.1
Supermemory 134.7 23.9 17.2 1.8 0.4 56.7
MemU 113.9 26.5 20.3 1.1 0.2 51.8
MemOS-1031 798.7 7.4 18.6 1.4 0.7 71.9
As shown in Table 5, MemOS not only achieved the best Personalized Response performance in both scenarios (with 0 turns and with 10 irrelevant turns) but also recorded the lowest preference unaware error. Importantly, it maintained acceptable context length control and low overall error rates. This demonstrates MemOS’s stable recognition of user preferences and robustness in long-term memory, highlighting its unique value in understanding user intent and providing highly personalized experiences.
Table 6: We evaluated our MemOS-1031 model against various baselines on the PersonaMem [103] benchmark, all built upon the same GPT-4o-mini[17] foundation LLM. We report the precision of the chosen one-in-four along with the context tokens used by each method. As the results show, MemOS-1031 achieved the best precision while maintaining acceptable context length control.
Method MIRIX Mem0 Zep Memobase MemU Supermemory MemOS-1031
@@
context tokens - 140 1657 2092 496 204 1424
Similarly, for the PersonaMem benchmark (Table 6), MemOS achieved the best precision while maintaining acceptable context length control, further validating its superior capability in handling dynamic user profiles and preferences.
6.3 Evaluation of chunk sizes and Top-K selection
@@
6.4 Evaluation of Memory Retrieval Robustness
Table 7: We evaluated the latency and LLM evaluation scores of various API methods on the LoCoMo benchmark under different QPS (queries per second) pressure of network request. Metrics reported include the P99, P90, and mean latency, as well as the success percentage rate for both add and search operations. Our results show that MemOS-1031 exhibited the most robustness, maintaining the lowest latency across nearly all metrics regardless of the QPS pressure. Furthermore, MemOS-1031 achieved a 100% success rate and maintained reasonable latency even under 100 QPS.
Add Search
QPS API P99(ms) $\downarrow$ P90(ms) $\downarrow$ mean latency (ms) $\downarrow$ success percentage (%) $\uparrow$ P99(ms) $\downarrow$ P90(ms) $\downarrow$ mean latency (ms) $\downarrow$ success percentage (%) $\uparrow$
10 qps Mem0 2841.0 1650.0 888.0 100.0 4637.0 1397.0 1089.0 100.0
Memobase 6169.1 1533.1 1126.2 99.9 14856.7 5974.7 2105.2 99.9
Supermemory 3467.4 1816.9 1446.3 100.0 2107.6 1457.4 1057.1 100.0
MemU 12070.6 7273.1 5077.9 64.2 63000.7 60539.5 47554.5 7.7
Zep 375.1 254.5 239.0 99.7 5348.7 614.8 571.7 99.9
MemOS-1031 376.4 211.6 191.9 100.0 777.1 528.4 440.5 100.0
40 qps Mem0 2624.0 1258.0 672.89 41.7 1723.0 1309.0 698.22 51.2
Memobase 7074.0 6251.5 2794.4 68.8 39985.3 18822.0 5515.5 97.8
Supermemory 60639.1 50951.5 24996.6 89.4 2050.8 1117.2 926.0 100.0
MemU 42129.9 35730.3 29033.0 7.5 61944.1 60565.2 41042.7 3.8
Zep 269.9 240.5 227.7 26.6 5802.5 1480.2 670.0 26.8
MemOS-1031 282.1 222 192.6 100.0 951.6 767.6 613.8 100.0
100 qps MemOS-1031 463.2 311.1 251.9 100.0 1171.1 872.5 741.2 100.0
We conduct a focused evaluation to analyze the efficiency and effectiveness of memory retrieval via network API. As shown in Table 7, we compare latency and success percentage for various API-based baselines under different Queries Per Second (QPS) request pressures. The baselines include Mem0, Memobase, Supermemory, MemU, and Zep, which are introduced in Section 6.1. The request contents are randomly sampled from the LoCoMo benchmark.
To test robustness, the metrics reported include the P99, P90, and mean latency, as well as the success percentage rate for both memory insertion (add) and retrieval (search) operations.
Our results show that MemOS exhibited the highest robustness with a 100% success percentage, maintaining the lowest latency across nearly all metrics regardless of the QPS pressure. Remarkably, MemOS achieved a 100% success rate and maintained reasonable latency even under 100 QPS. This demonstrates that MemOS’s hybrid semantic organization and activation-based memory loading can achieve superior and highly stable performance across a wide range of QPS pressure.